# Quadratic dispersion relations in gapless frustration-free systems

### v1

@@ Abstract
Recent case-by-case studies revealed that the dispersion of low energy excitations in gapless frustration-free Hamiltonians is often quadratic or softer. In this work, we argue that this is actually a general property of such systems. By combining a previous study by Bravyi and Gosset and the min-max principle, we prove this hypothesis for models with local Hilbert spaces of dimension two that contains only nearest-neighbor interactions on cubic lattice. This may be understood as a no-go theorem realizing gapless phases with linearly dispersive excitations in frustration-free Hamiltonians. We also provide examples of frustration-free Hamiltonians in which the plane-wave state of a single spin flip does not constitute low energy excitations.
@@ I Introduction
Accurately investigating the ground states and excited states of quantum many-body systems is often challenging. It is hence common practice to analyze the system of interest using the simplest possible model that belongs to the same phase. Frustration-free systems, in particular, are relatively easy to handle, and their exact ground states and excited states are sometimes accessible. For these reasons, frustration-free systems have been widely used as representative models of various phases of matter. Therefore, elucidating the general properties of frustration-free systems and understanding the limits of the phases that can be realized by frustration-free Hamiltonians hold significant importance.
@@
Numerous frustration-free systems posses an excitation gap. Notable examples include the Majumdar–Ghosh (MG) model, known as a toy model realizing spontaneous breaking of discrete translational symmetry, the Affleck–Kennedy–Lieb–Tasaki (AKLT) model, representing a symmetry-protected topological phase, and the Kitaev Toric code model, serving as the canonical model for topologically ordered phases Auerbach (2012); Tasaki (2020). The parent Hamiltonians of Matrix Product States (MPS) also feature an excitation gap Fannes et al. (1992).
@@
There are also frustration-free Hamiltonians with gapless excitations. For example, the uncle Hamiltonian of MPS Fernández-González et al. (2015), various models at critical points, and models related to spontaneously broken continuous symmetries Ogunnaike et al. (2023); Watanabe et al. (2023) are gapless. Through recent investigation of these models, the general properties of frustration-free systems have gradually become clearer.
@@
In this paper, we discuss two conjectures regarding frustration-free systems. The first is about the finite-size gap between degenerate ground states. Generally, even if the ground states are degenerate in the thermodynamic limit, their energy eigenvalues in a finite-size system do not precisely match due to the finite-size gap. However, in frustration-free systems, the ground states are exactly degenerate even before taking the thermodynamic limit. In other words, we can judge whether a frusration-free system is gapped or not by looking at the limiting behavior of the difference between the ground state energy and the second lowest energy level.
@@
The second conjecture concerns the dispersion relation of low-energy excitations in frustration-free systems. Although gapless modes typically have a linear dispersion, it is known that the ferromagnetic Heisenberg model, which is a representative frustration-free Hamiltonian, have a spin wave excitation with a quadratic dispersion Auerbach (2012). In fact, our second conjecture asserts that gapless frustration-free systems generally have a quadratic or softer dispersion relation. Although there are several arguments supporting this statement in literature (for example, see Refs. Palle and Benton, 2021; Ogunnaike et al., 2023; Watanabe et al., 2023; Ren et al., 2024), they are applicable only to the case where the low-energy excitations are obtained by applying a sum of local operators to a ground state. We provide a general proof for spin-1/2 models with nearest-neighbor interaction on cubic lattice based on the result of Bravyi and Gosset Bravyi and Gosset (2015). We also discuss examples with longer-range interactions and show that gapless excitations with quadratic dispersion may not be obtained by a sum of local operators.
@@ II Overview
II.0.1 Setting and definitions
In this work, we consider Hamiltonians defined on a finite $d$-dimensional lattice $\Lambda$:
$\hat{H}=\sum_{\bm{r}\in\Lambda}\hat{H}_{\bm{r}}.$ (1)
The local Hilbert space on each lattice site $\bm{r}\in\Lambda$ is assumed to be finite dimensional, which is denoted by $D_{0}$. Let $|\Lambda|$ be the number of lattice sites in $\Lambda$. Then the dimension of the total Hilbert space is $D=D_{0}^{|\Lambda|}$. We assume that the Hamiltonian is local in the sense that $\hat{H}_{\bm{r}}$ acts nontrivially on the sites within a finite distance $R$ from $\bm{r}$.
@@ II.0.1 Setting and definitions
The thermodynamic limit $|\Lambda|\to\infty$ is taken by a sequence of increasing system size.
The system is said to be gapped if there exists an integer $N_{\mathrm{deg}}$ ($1\leq N_{\mathrm{deg}}<D$), which may depend on the system size, such that
$\lim_{|\Lambda|\to\infty}\big{(}E_{N_{\mathrm{deg}}}-E_{1}\big{)}=0,$ (3)
$\Delta\coloneqq\lim_{|\Lambda|\to\infty}\big{(}E_{N_{\mathrm{deg}}+1}-E_{N_{\mathrm{deg}}}\big{)}\neq 0.$ (4)
@@
Otherwise the system is gapless. The limit in Eq. (3) defines the excitation gap $\Delta$. When $N_{\mathrm{deg}}\geq 2$, the quantity $E_{N_{\mathrm{deg}}}-E_{1}$ is referred to as the finite-size gap among degenerate ground states. When the degeneracy originates from spontaneous breaking of a discrete symmetry or formation of a topological order, it generically decays exponentially with the system size. Even when $N_{\mathrm{deg}}$ converges to a finite number in the large $|\Lambda|$ limit, its value for a finite system may be ambiguous due to the possible system-size dependence.
The Hamiltonian is translation invariant if there exists an operator $\hat{T}_{\bm{a}}$ for each lattice vector $\bm{a}$ such that $\hat{T}_{\bm{a}}\hat{H}_{\bm{r}}=\hat{H}_{\bm{r}+\bm{a}}\hat{T}_{\bm{a}}$ for all $\bm{r}\in\Lambda$. This is possible only when the periodic boundary condition (PBC) is imposed.
@@
The Hamiltonian is frustration-free if a ground state $|{\Phi_{0}}\rangle$ of $\hat{H}$ is a simultaneous ground state of every local Hamiltonian $\hat{H}_{\bm{r}}$ Tasaki (2020). In this case, without loss of generality and without modifying the ground states and low-energy excitations of the system, one may assume that each $\hat{H}_{\bm{r}}$ is a projector, i.e., $\hat{Q}_{\bm{r}}^{2}=\hat{Q}_{\bm{r}}$ and $\hat{Q}_{\bm{r}}|{\Phi_{0}}\rangle=0$ for all $\bm{r}\in\Lambda$. Under this choice, the ground state energy becomes $E_{1}=0$.
II.0.2 Conjectures on frustration-free Hamiltonian
Let us state our conjectures one by one. The first conjecture is that, when the Hamiltonian is frustration-free and gapped, the finite-size gap among ground states precisely vanishes before taking the thermodynamic limit:
$E_{N_{\mathrm{deg}}}=E_{1}.$ (5)
For example, in the AKLT model for spin $s=1$ chain under OBC, the four-fold ground state degeneracy due to the edge modes is exact. The four-fold topological degeneracy in the Kitaev toric code under PBC is also exact.
This implies the following criterion for the excitation gap in frustration-free systems. Let $\tilde{N}_{\mathrm{deg}}$ ($1\leq\tilde{N}_{\mathrm{deg}}<D$) be the number of exact zero energy states, i.e.,
$E_{\tilde{N}_{\mathrm{deg}}}=E_{1},\quad E_{\tilde{N}_{\mathrm{deg}}+1}\neq E_{1},$ (6)
which may also depend on the system size (see, for example, Eq. (45) below). The Hamiltonian is gapped if and only if
$\lim_{|\Lambda|\to\infty}\big{(}E_{\tilde{N}_{\mathrm{deg}}+1}-E_{1}\big{)}>0.$ (7)
@@ II.0.2 Conjectures on frustration-free Hamiltonian
Namely, one can set $N_{\mathrm{deg}}=\tilde{N}_{\mathrm{deg}}$. Actually this criterion has been used as the definition of the excitation gap in previous studies (for example, see Refs. Bravyi and Gosset, 2015; Gosset and Mozgunov, 2016; Gosset and Huang, 2016). Note that $\tilde{N}_{\mathrm{deg}}$ is well-defined even for gapless systems unlike $N_{\mathrm{deg}}$.
@@
The second conjecture is that, when the Hamiltonian is translation-invariant, frustration-free, and gapless, there exists a family of variational states $|\Psi_{\bm{k}}\rangle$ that are orthogonal to all the ground states and satisfy
$\hat{T}_{\bm{a}}|\Psi_{\bm{k}}\rangle=e^{-i\bm{k}\cdot\bm{a}}|\Psi_{\bm{k}}\rangle,$ (8)
@@
for some $\bm{k}_{0}$. Namely, the low energy excitation of the system is not linearly dispersive about the gapless point at $\bm{k}=\bm{k}_{0}$.
For systems under OBC, the result of Refs. Knabe, 1988; Anshu, 2020; Gosset and Mozgunov, 2016 gives us the bound $E_{\tilde{N}_{\mathrm{deg}}+1}-E_{1}\leq O(L^{-2})$. Although this is consistent with our conjecture, it is not directly applicable to the PBC case Watanabe et al. (2023).
II.0.3 Min-max principle
Next, let us review the min-max principle, which supports our conjectures. We compare two Hamiltonians $\hat{H}$ and $\hat{H}^{\prime}$ acting on a $D$-dimensional Hilbert space. The eigenvalues of $\hat{H}$ and $\hat{H}^{\prime}$ are, respectively, denoted by $E_{j}$ and $E_{j}^{\prime}$ ($j=1,2,\cdots,D$) in the increasing order. Suppose that $\hat{V}\coloneqq\hat{H}^{\prime}-\hat{H}$ is positive semi-definite:
$\hat{V}\geq 0.$ (10)
That is, $\langle\Phi|\hat{V}|\Phi\rangle\geq 0$ for any state $|\Phi\rangle$. In other words, all the eigenvalues of $\hat{V}$ are nonnegative. In this setting, the min-max principle states that
$E_{j}^{\prime}\geq E_{j}$ (11)
for every $j=1,2,\cdots,D$ (see Theorem A.7 of Ref. Tasaki (2020)). This statement is trivial when the two Hamiltonians can be diagonalized by a common unitary operator, but it holds more generally. Since the proof is elementary, we review it in Appendix A.
Two corollaries follow immediately from the min-max principle. Let us consider two frustration free Hamiltonians $\hat{H}=\sum_{\bm{r}\in\Lambda}\hat{Q}_{\bm{r}}$ and $\hat{H}^{\prime}=\sum_{\bm{r}\in\Lambda}\hat{Q}_{\bm{r}}^{\prime}$. Suppose that $\hat{V}=\hat{H}^{\prime}-\hat{H}$ is positive-semidefinite. Then the inequality in (11) implies that the number of zero-energy states of $\hat{H}$ cannot be smaller than that of $\hat{H}^{\prime}$:
$\tilde{N}_{\mathrm{deg}}\geq\tilde{N}_{\mathrm{deg}}^{\prime}.$ (12)
Furthermore,
$\lim_{|\Lambda|\to\infty}E_{j}^{\prime}=0\quad\Rightarrow\quad\lim_{|\Lambda|\to\infty}E_{j}=0.$ (13)
This relation suggests that if (i) $\hat{H}$ is gapped and (ii) $\tilde{N}_{\mathrm{deg}}$ is bounded by a system-size independent constant, then $\hat{H}^{\prime}$ is also gapped. This relation was previously used for interacting Kitaev chains in Refs. Katsura et al., 2015; Wouters et al., 2018.
III 1D Examples with nearest-neighbor interactions
@@ II.0.3 Min-max principle
$\hat{H}^{\mathrm{pbc}}\coloneqq\sum_{x=1}^{L}\hat{Q}_{x,x+1},\quad\hat{H}^{\mathrm{obc}}\coloneqq\sum_{x=1}^{L-1}\hat{Q}_{x,x+1}.$ (14)
Here, $\hat{Q}_{x,x+1}=\hat{Q}_{x,x+1}^{2}$ is a projector nontrivially acting on the spins at $x$ and $x+1$. The interaction range of this Hamiltonian is 1. The number of $+1$ eigenvalues of a projector is called the rank. The Hamiltonian $\hat{H}^{\mathrm{pbc}}$ has the translation symmetry $\hat{T}$ which satisfies $\hat{T}\hat{Q}_{x-1,x}=\hat{Q}_{x,x+1}\hat{T}$.
The seminal work by Bravyi and Gosset Bravyi and Gosset (2015) showed that $\hat{H}^{\mathrm{obc}}$ can be gapless only when the projector $\hat{Q}_{x,x+1}$ is unitary equivalent to Eqs. (19), (20) for the rank 1 case and Eq. (38) for the rank 2 case. Then the min-max principle implies the same for $\hat{H}^{\mathrm{pbc}}$ at least when $\tilde{N}_{\mathrm{deg}}$ under OBC is finite, since $\hat{H}^{\mathrm{pbc}}-\hat{H}^{\mathrm{obc}}=\hat{Q}_{L,1}\geq 0$ is positive semi-definite. We will see that the fully polarized state
$|{\Phi_{0}}\rangle\coloneqq\bigotimes_{x=1}^{L}|0\rangle_{x}=|0\cdots 0\rangle$ (15)
is a common ground state of these Hamiltonians. Furthermore, the plane-wave state of a single spin flip
$|{\Psi_{k}}\rangle\coloneqq\frac{1}{\sqrt{L}}\sum_{x=1}^{L}e^{ikx}|x\rangle,$ (16)
$|x\rangle\coloneqq\hat{s}_{x}^{-}|{\Phi_{0}}\rangle=|\underbrace{0\cdots 0}_{x-1}1\underbrace{0\cdots 0}_{L-x}\rangle$ (17)
with $k\coloneqq 2\pi m/L$ ($m=1,2,\cdots L$) is a common variational state among these examples and has a quadratic dispersion relation. That is,
$E(k)$ $=\langle\Psi_{k}|\hat{H}^{\mathrm{pbc}}|{\Psi_{k}}\rangle=\sum_{x=1}^{L}\big{\|}\hat{Q}_{x,x+1}|{\Psi_{k}}\rangle\big{\|}^{2}$
$=\frac{1}{L}\sum_{x,x^{\prime}=1}^{L}e^{ik(x-x^{\prime})}\langle x^{\prime}|\hat{H}^{\mathrm{pbc}}|x\rangle=O\big{(}(k-k_{0})^{2}\big{)}$ (18)
for some $k_{0}$.

### v2

@@ Abstract
Recent case-by-case studies revealed that the dispersion of low energy excitations in gapless frustration-free Hamiltonians is often quadratic or softer. In this worl, we argue that this is actually a general property of such systems. By combining the previous study by Bravyi and Gosset and the min-max principle, we prove this hypothesis for models with local Hilbert spaces of dimension two that contains only nearest-neighbor interactions on cubic lattice. This may be understood as a no-go theorem realizing gapless phases with linearly dispersive excitations in frustration-free Hamiltonians. We also provide examples of frustration-free Hamiltonians in which the plane-wave state of a single spin flip does not constitute low energy excitations.
@@ I Introduction
Accurately investigating the ground states and excited states of quantum many-body systems is often challenging. It is hence common practice to analyze the system of interest using the simplest possible model that belongs to the same phase. Frustration-free (FF) systems, in particular, are relatively easy to handle, and their exact ground states and excited states are sometimes accessible. For these reasons, FF systems have been widely used as representative models of various phases of matter. Therefore, elucidating the general properties of FF systems and understanding the limits of the phases that can be realized by FF Hamiltonians hold significant importance.
@@
Numerous FF systems possess an excitation gap. Notable examples include the Majumdar–Ghosh (MG) model, known as a toy model realizing spontaneous breaking of discrete translational symmetry, the Affleck–Kennedy–Lieb–Tasaki model, representing a symmetry-protected topological phase, and the Kitaev toric code model, serving as the canonical model for topologically ordered phases Auerbach (2012); Tasaki (2020). The parent Hamiltonians of Matrix Product States (MPS) also feature an excitation gap Fannes et al. (1992).
@@
There are also FF Hamiltonians with gapless excitations. For example, the uncle Hamiltonian of MPS Fernández-González et al. (2015), various models at critical points, and models related to spontaneously broken continuous symmetries Ogunnaike et al. (2023); Watanabe et al. (2023) are gapless. Through their recent investigations, the general properties of FF systems have gradually become clearer.
@@
In this worl, we discuss two conjectures regarding FF systems. The first one is about the finite-size splitting between degenerate ground states. Generally, even if the ground states are degenerate in the thermodynamic limit, their energy eigenvalues in a finite system do not match due to the finite-size splitting. However, in FF systems, the ground states are exactly degenerate even before taking the thermodynamic limit. In other words, whether a frusration-free system is gapped or not can be judged by looking at the limiting behavior of the difference between the ground state energy and the second lowest energy level.
@@
The second conjecture concerns the dispersion relation of low-energy excitations in FF systems. Although gapless modes typically have a linear dispersion, it is known that the ferromagnetic Heisenberg model, a representative FF Hamiltonian, have a spin wave excitation with a quadratic dispersion $E(k)=2\sin^{2}(k/2)$ Auerbach (2012). In fact, our second conjecture asserts that gapless FF systems always have a quadratic or softer dispersion relation. Although there are several arguments supporting this statement in literature Palle and Benton (2021); Ogunnaike et al. (2023); Watanabe et al. (2023); Ren et al. (2024), they are applicable only to the case where the low-energy excitations are obtained by applying a sum of local operators to a ground state. We provide a general proof for spin-1/2 models with nearest-neighbor interaction on cubic lattice based on Ref. Bravyi and Gosset, 2015. We also discuss examples with longer-range interactions and show that gapless excitations with quadratic dispersion may not be obtained by a sum of local operators.
@@ II Overview
II.1 Setting and definitions
In this work, we consider Hamiltonians
$\hat{H}=\sum_{\bm{r}\in\Lambda}\hat{H}_{\bm{r}}$ (1)
defined on a finite $d$-dimensional lattice $\Lambda$. The local Hilbert space on each lattice site $\bm{r}\in\Lambda$ is finite dimensional, which is denoted by $D_{0}$. Let $|\Lambda|$ be the number of lattice sites in $\Lambda$. Then the dimension of the total Hilbert space is $D=D_{0}^{|\Lambda|}$. We assume that the Hamiltonian is local in the sense that $\hat{H}_{\bm{r}}$ acts nontrivially on the sites within a finite distance $R$ from $\bm{r}$. The Hamiltonian is translation invariant if $\hat{T}_{\bm{a}}\hat{H}_{\bm{r}}=\hat{H}_{\bm{r}+\bm{a}}\hat{T}_{\bm{a}}$ for each lattice vector $\bm{a}$ and $\bm{r}\in\Lambda$. This is possible only when the periodic boundary condition (PBC) is imposed.
@@ II.1 Setting and definitions
The thermodynamic limit $|\Lambda|\to\infty$ is taken by a sequence of increasing system size. The system is said to be gapped if there exists an integer $N_{\mathrm{deg}}$ ($1\leq N_{\mathrm{deg}}<D$), which may depend on the system size, such that $\lim_{|\Lambda|\to\infty}\big{(}E_{N_{\mathrm{deg}}}-E_{1}\big{)}=0$ and $\lim_{|\Lambda|\to\infty}\big{(}E_{N_{\mathrm{deg}}+1}-E_{N_{\mathrm{deg}}}\big{)}\neq 0$. Otherwise the system is gapless, and the spectrum above the ground state is continuous. When $N_{\mathrm{deg}}\geq 2$, the quantity $E_{N_{\mathrm{deg}}}-E_{1}$ is referred to as the finite-size splitting among degenerate ground states. When the degeneracy originates from spontaneous breaking of a discrete symmetry or formation of a topological order, it generically decays exponentially with the system size. Even when $N_{\mathrm{deg}}$ converges to a finite number in the large $|\Lambda|$ limit, its value for a finite system may be ambiguous due to the possible system-size dependence.
@@
The Hamiltonian is FF if a ground state $|{\Phi_{0}}\rangle$ of $\hat{H}$ is a simultaneous ground state of every local Hamiltonian $\hat{H}_{\bm{r}}$ Tasaki (2020). In this case, without loss of generality nor modifying the ground states and low-energy excitations of the system, one may assume that each $\hat{H}_{\bm{r}}$ is a projector, i.e., $\hat{Q}_{\bm{r}}^{2}=\hat{Q}_{\bm{r}}$ and $\hat{Q}_{\bm{r}}|{\Phi_{0}}\rangle=0$ for all $\bm{r}\in\Lambda$. Under this choice, the ground state energy becomes $E_{1}=0$.
II.2 Conjectures
Let us state our conjectures one by one.
II.2.1 Conjecture 1
The first conjecture is that, when the Hamiltonian is FF and gapped, the finite-size splitting among ground states precisely vanishes before taking the thermodynamic limit; i.e., $E_{N_{\mathrm{deg}}}=E_{1}$. For example, in the AKLT model for spin $s=1$ chain under OBC, the four-fold ground state degeneracy due to the edge modes is exact. The four-fold topological degeneracy in the Kitaev toric code under PBC is also exact. This observation implies the following criterion for the excitation gap in FF systems. Let $\tilde{N}_{\mathrm{deg}}$ ($1\leq\tilde{N}_{\mathrm{deg}}<D$) be the number of exact zero energy states, i.e., $E_{\tilde{N}_{\mathrm{deg}}}=0$ and $E_{\tilde{N}_{\mathrm{deg}}+1}\neq 0$, which may also depend on the system size. The Hamiltonian is gapless if and only if
$\lim_{|\Lambda|\to\infty}E_{\tilde{N}_{\mathrm{deg}}+1}=0.$ (3)
@@ II.2.1 Conjecture 1
Actually this criterion has been used as the definition of the excitation gap in previous studies Bravyi and Gosset (2015); Gosset and Mozgunov (2016); Gosset and Huang (2016). Note that $\tilde{N}_{\mathrm{deg}}$ is well-defined even for gapless systems unlike $N_{\mathrm{deg}}$.
II.2.2 Conjecture 2
@@
The second conjecture is that, when the Hamiltonian is translation-invariant, FF, and gapless, there exists a family of variational states $|\Psi_{\bm{k}}\rangle$ that are orthogonal to all the ground states, are eigenstates of translation operator $\hat{T}_{\bm{a}}|\Psi_{\bm{k}}\rangle=e^{-i\bm{k}\cdot\bm{a}}|\Psi_{\bm{k}}\rangle$, and have a quadratic dispersion about the gapless point at $\bm{k}=\bm{k}_{0}$:
@@
For systems under OBC, the result of Refs. Knabe, 1988; Anshu, 2020; Gosset and Mozgunov, 2016; Lemm and Xiang, 2022 gives us the bound $E_{\tilde{N}_{\mathrm{deg}}+1}=O(L^{-2})$. Although this is consistent with our conjecture, it is not directly applicable to the PBC case Watanabe et al. (2023).
In Sec. III, we give a proof of this conjecture for $s=1/2$ spin chain with nearest neighbor interactions. We further generalize the results to higher dimensions in Sec. IV.
II.3 Min-max principle
Our discussions below are based on a mathematical theorem called the min-max principle (see, theorem A.7 of Ref. Tasaki, 2020). We compare two Hamiltonians $\hat{H}$ and $\hat{H}^{\prime}$ acting on the same $D$-dimensional Hilbert space. The eigenvalues of $\hat{H}$ and $\hat{H}^{\prime}$ are, respectively, denoted by $E_{j}$ and $E_{j}^{\prime}$ ($j=1,2,\cdots,D$) in the increasing order. Suppose that $\hat{V}\coloneqq\hat{H}^{\prime}-\hat{H}$ is positive semi-definite $\hat{V}\geq 0$. That is, $\langle\Phi|\hat{V}|\Phi\rangle\geq 0$ for any state $|\Phi\rangle$. In this setting, the min-max principle states that
$E_{j}^{\prime}\geq E_{j}\quad({}^{\forall}j=1,2,\cdots,D)$ (5)
For readers’ convenience, we review the proof in Appendix A. This is trivial when the two Hamiltonians can be diagonalized by a common unitary operator, but it holds more generally.
Two corollaries follow immediately from the theorem. Let us consider two FF Hamiltonians $\hat{H}=\sum_{\bm{r}\in\Lambda}\hat{Q}_{\bm{r}}$ and $\hat{H}^{\prime}=\sum_{\bm{r}\in\Lambda}\hat{Q}_{\bm{r}}^{\prime}$. Suppose that $\hat{V}=\hat{H}^{\prime}-\hat{H}$ is positive-semidefinite. Then the minimax principle implies that the number of zero-energy states of $\hat{H}$ cannot be smaller than that of $\hat{H}^{\prime}$: $\tilde{N}_{\mathrm{deg}}\geq\tilde{N}_{\mathrm{deg}}^{\prime}$. Furthermore, if (i) $\hat{H}$ is gapped and (ii) $\tilde{N}_{\mathrm{deg}}$ is bounded by a system-size independent constant, then $\hat{H}^{\prime}$ is also gapped. This is because, if $\hat{H}^{\prime}$ were gapless, an increasing number of eigenstates of $\hat{H}^{\prime}$ have lower energy than $E_{\tilde{N}_{\mathrm{deg}}+1}>0$ and eventually $E_{\tilde{N}_{\mathrm{deg}}+1}>E_{\tilde{N}_{\mathrm{deg}}+1}^{\prime}$ holds for a sufficiently large $L$. This relation was previously used for interacting Kitaev chains Katsura et al. (2015); Wouters et al. (2018).
III 1D system with nearest-neighbor interactions
@@ II.3 Min-max principle
Figure 1: The illustration of the 2D model and its decomposition into 1D models.
The seminal work by Bravyi and Gosset Bravyi and Gosset (2015) showed that the Hamiltonian $\hat{H}^{\mathrm{obc}}\coloneqq\sum_{x=1}^{L-1}\hat{Q}_{x,x+1}$ under open boundary condition (OBC) can be gapless only when the projector $\hat{Q}_{x,x+1}$ is unitary equivalent to Eqs. (7), (8) for the rank 1 case and Eq. (13) for the rank 2 case. Then the min-max principle implies the same for $\hat{H}$ under PBC, since $\hat{H}-\hat{H}^{\mathrm{obc}}=\hat{Q}_{L,1}\geq 0$ is positive semi-definite. We will see that the fully polarized state $|{\Phi_{0}}\rangle\coloneqq\bigotimes_{x=1}^{L}|0\rangle_{x}=|0\cdots 0\rangle$ is a common ground state of these Hamiltonians. Furthermore, the plane-wave state
$|{\Psi_{k}}\rangle\coloneqq\frac{1}{\sqrt{L}}\sum_{x=1}^{L}e^{ikx}\hat{s}_{x}^{-}|{\Phi_{0}}\rangle$ (6)
with $k\coloneqq 2\pi m/L$ ($m=0,1,2,\cdots L-1$) is a common variational state with a quadratic dispersion relation.